OpenStudy (anonymous):

3x^2 – 5x – 2 Factor problem and show the work.

OpenStudy (anonymous):

you'll have to use delta and all that stuff

OpenStudy (anonymous):

\[\Delta= b^2 - 4ac\] a=3 b= -5 c = -2 \[\Delta = (-5)^2 - 4(3)(-2) = 49\] so : \[ x = -b \pm \sqrt{\Delta}/2a\] substitute delta and do the calculation :) give it a try.

OpenStudy (anonymous):

what will your answer be?

OpenStudy (anonymous):

Your answer should be:\[x1 = 2 \] and \[x2 = -1/3\] did you get that?

OpenStudy (anonymous):

nice sstarica you are smart

OpenStudy (anonymous):

thank you ^_^

OpenStudy (anonymous):

If the directions said to factor, then you can't use the quadtratic formula. The answer would be (3x + 1)(x - 2) there is no solution if you don't have an = sign to begin with.

OpenStudy (anonymous):

u also can use the calculator ..its very easy and efficient..

OpenStudy (anonymous):

but you've got to learn the way too :) never trust techonology in the end, they fail you.